A package is dropped from a helicopter flying at a height of 7.5 m while traveling 15 m/s. how far does the package travel horizontally before it hits the ground below?


1
Expert's answer
2022-05-19T05:43:00-0400

"L=v_0\sqrt{2H/g}=15*\sqrt{2*7.5/9.8}=19\:\rm m"